OpenStudy (anonymous):

Varia is studying abroad in Europe. She is required pay $3,500 (in US dollars) per year to the university, however, she must pay in euros. How many euros can Varia expect to pay per month to the university? (Round to the nearest whole euro.) 1.3687 euros = 1 US dollar euros

OpenStudy (anonymous):

Cost of 1 year = $3,500 Cost of 1 month is given by \[\frac{ 3500 }{ 12 }=\frac{ 875 }{ 3 }=291.667\] This cost is in dollars so convert it to euros 1$=1.3687eu 291.667$=291.667*1.3687eu \[291.667\times1.3687=399.2\in\]

OpenStudy (anonymous):

399euro

OpenStudy (anonymous):

thankyou